OpenStudy (anonymous):

Find the equation of the tangent line to the curve at the point (3, 2). x^2+xy+2y^2 = 23

OpenStudy (zarkon):

\[\frac{dy}{dx}=\frac{-(2x+y)}{x+4y}\] evaluate at (3,2) plug into \[y-y_1=m(x-x_1)\]

OpenStudy (anonymous):

plz help me w/ anotehr one..ill psot it